2c). Solve each of the following


I) x^2=72


II) (3n-1)^2=26


1
Expert's answer
2021-10-01T09:50:36-0400

i) "x^2=72\to x=\sqrt{72}, x=-\sqrt{72}."


ii) "(3n-1)^2=26\to 3n-1=\sqrt{26}\to""n=\frac{\sqrt{26}+1}{3}, n=-\frac{\sqrt{26}+1}{3}" .